# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v2

@@ B.3 Dependence on Latent Layer of Study
Figure 14: This network shows the PPP method applied over networks of 333 hidden layer block, 181818 neurons and using input-output normalisation as described in Apps. A and D. PPP was applied to the $0^{\text{th}}$ hidden layer. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in identical fashion to Fig. 1. One can observe that representations are absent in anisotropic rows 1, 2, and 4, likely a result of anti-alignment, whereas some axis-aligned structure is present in rows 3 and 5. Isotropic rows demonstrate larger, more continuous representational structures, but also demonstrate clustering such that they are not distributed smoothly over the representation space.
@@
Overall, in many networks tested, anisotropic structure appeared progressively clearer and denser in later latent layers, whilst isotropic networks tended to produce progressively more continuous distributions in later latent layers. This seemed to be a trend across many of the networks studied, but sporadic network-to-network differences also occurred. This may be due to the increased number of an/isotropies in the forward pass or indirectly through the backwards pass, yet further study may elucidate the exact reason why progressively clearer structure may emerge in later layers.
@@ B.4 Results for Leaky-ReLU
Figure 18: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 333 hidden layer blocks and width of 181818 neurons per block. The top five rows consist of the autoencoder using the standard Leaky-ReLU, while the bottom five rows consist of the same autoencoders using the isotropic Leaky-ReLU. All five of the top five rows demonstrate clear anti-aligned quantisation in representations, predominantly in a single anti-aligned axis. These emerge early in training and become progressively more (approximately) discrete. This contrasts with the isotropic Leaky-ReLU examples, which approach a more continuous, evenly distributed embedding distribution, consistent with previous observations. Additionally, at relatively smaller magnitudes, the Leaky-ReLU results consistently show some deviation from the discrete-like rays; this remains unexplained.
Figure 19: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 222 hidden layer blocks and width of 181818 neurons per block. The top five rows consist of the autoencoder using the standard Leaky-ReLU, while the bottom five rows consist of the same autoencoders using the isotropic Leaky-ReLU. These results are highly consistent with Fig. 18, yet the 3rd row also shows the addition of axis-aligned structure.
@@
In particular, the $0^{\text{th}}$ layer standard Leaky-ReLU results always produced two narrow beams arranged antipodally, along a single diagonal. This anti-aligned structure occurs in most cases, and the standard Leaky-ReLU only rarely displays an axis-aligned structure. This single-diagonal may be indicative of the $S_{n}$ group, which does not include sign flips, unlike tanh’s $B_{n}$ group. Therefore, only a single diagonal may appear populated, as opposed to other diagonal directions, which could result from the various mirrors applicable in the $B_{n}$ group. This nuanced difference between demonstrations appears to closely follow from the differences between symmetry groups. These differing representations can be considered to support the influence of different symmetry groups on representations, with nuances following in a taxonomic fashion from the group structures. Particularly, the orthants containing ±$\pm\vec{1}$ are the only non-degenerate orthants produced by the $S_{n}$ symmetry, and the representations appear to be aligning with these, suggesting the remaining orthants appear not to produce computationally useful maps.
@@ B.5 MNIST Results
An initial example is displayed in Fig. 22 and 23 — the latter showing the same plot, with colour levels adjusted to make faint structure more observable for analysis. Both are included for comparability and transparency.
Figure 22: Displays the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 333-hidden-layer blocks consisting of 181818 neurons per block. This architecture is comparable to those in App. D in every way except for the dimensions of CIFAR, which is exchanged for MNIST shape on input-output layers. Overall, one can observe that isotropic networks in the bottom five rows continue to produce smooth distributions that fill the representation space. In comparison, anisotropic networks on the top five rows tend to have many representations centred around the origin, and some produce axis-aligned features.
@@
Several of the $0^{\text{th}}$ layer plots show a closer grouping of representations about the origin. This is believed to be due to many of the MNIST examples’ components reaching the extreme values for the space, 0 and 111, with fewer intermediate values. Often, this occurs repeatedly in specific patches due to the overlapping morphologies of MNIST numbers, which may produce clustered representations inherent to the dataset. This makes the MNIST dataset more anisotropic, featuring already dense, discrete clusters about the corners of an embedded hyper-cube — significantly different from the more intermediate RGB values arising from CIFAR. Slight band-groupings of representations are also sometimes observable, likely due to the same representation being projected in multiple planes, as well as the precision of the dataset, which produces slight clusterings. However, MNIST’s greater inherent anisotropy appears not to significantly alter the results from CIFAR, likely due to the layer bottleneck embedding these higher-dimensional representations more evenly in the latent layer during initialisation.
@@
Figure 24: Displays the PPP method applied to the $1^{\text{st}}$ latent layer of an autoencoder with 333-hidden-layer blocks consisting of 181818 neurons per block. Row-wise, these are the exact same networks as displayed in Fig. 22 and 23, but for the next latent layer. One can observe an axis-aligned structure in the top five rows, corresponding to anisotropic standard-tanh, and this is consistent with all previous experiments. In contrast, a smoother, approximately continuous distribution is observed for the bottom five rows, which are isotropic, and is consistent with this trend.
@@ Appendix C Analysis of Performance and Pathological Indications
Figure 26: These plots show the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green) reconstruction errors on test-set CIFAR against the number of training epochs. The networks have a width of 181818 neurons, and their depth increases to the right. In the leftmost plot, unnormalised isotropy has the lowest error by a very slight margin. The central two plots show that normalised isotropic networks have the lowest error by a large margin. The rightmost plot again shows unnormalised isotropy with the lowest error and a larger margin. Overall, a trend is not apparent, but isotropy tends to outperform by a greater margin.
@@
An instance of 323232 neurons was also computed, and this example did show unnormalised-anisotropic functions performing better as shown in Fig. 27. However, these were not the best-performing models overall.
Figure 27: This plot shows the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green), reconstruction error on test-set CIFAR against the number of training epochs. The networks have a width of 323232 neurons, and 222 hidden layer blocks. In this counterexample, the unnormalised-anisotropic network had the lowest reconstruction error.
@@
Figure 28: This plot shows the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) reconstruction errors for test-set MNIST against the number of training epochs, network width increases rightward over columns, and depth increases downward over rows (flipped ordering from Fig. 25). It can be observed that reconstruction error is lower in wider and less-deep networks. Isotropic networks outperform anisotropic networks in every case, except for the 333-hidden-layer block and 323232 neurons, where anisotropic networks exhibit a slight marginal success. The best performing network is the isotropic network with 0 hidden layer blocks and 323232 neurons — this network is also thought to have the greatest representation capacity.
@@
Figure 29: These plots show the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green), reconstruction errors on test-set MNIST against the number of training epochs. The networks have a width of 181818 neurons, and a depth of 0 hidden layer blocks. It can be observed that all networks tend to a similar reconstruction error, except for the normalised-isotropy graph, which reaches a significantly lower reconstruction error.
@@
Figure 30: These plots show the Leaky-ReLU-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green) reconstruction errors on test-set CIFAR against the number of training epochs. The networks have a width of 181818 neurons, and their depth increases to the right. In all instances, normalised or unnormalised isotropic-Leaky-ReLU reaches the lowest reconstruction error by a large margin. Anisotropic networks also appear to exhibit significantly greater noise and sample-to-sample variation in test-set accuracies across the differing training epochs.
@@ Appendix D Further Details: Network Architectures and Training
[…]
Finally, following from App. A, the MNIST and CIFAR datasets are sometimes normalised over their training sets to reduce statistical geometry artefacts. This involves an element-wise calculation of the mean tensor over all training samples, as well as for the standard deviation statistic. Where the standard deviation is found to be 0, those elements are replaced with the value 111. When normalisation is used, every ingoing sample per batch is standardised by subtracting the mean and dividing by the standard deviation, where the non-zero standard deviation is made equal to one.
Throughout all experiments, a batch size of 242424 was maintained, alongside a learning rate for PyTorch’s Paszke et al. [2019] momentum gradient descent, with momentum 0.90.90.9, a learning rate 0.080.080.08, and using a mean-squared error loss for reconstruction. Parameter initialisation was the standard Xavier-Normal Glorot and Bengio [2010] for weights and zero-initialised for biases, this is due to their left-, right- probabilistic invariances Bird [2025a]. However, standard multivariate normal or orthogonal initialisation would have been equally valid choices. Still, Xavier-Normal was deemed more conventional, though its gain is only optimised for anisotropic activation functions Bird [2025a]. All values were kept in float64.
[…]
Data was gathered using combination-PPP, with $\epsilon=0.75$, which was found to be a good overall value. Plots were compiled using Gaussians centred at each representation’s projected coordinates, and the colour map indicates representation density per region, with a maximum value of the mean plus five standard deviations and a minimum of zero. Additionally, in the anisotropic cases, so many representations became aligned with the axes that computing the Gaussian maps became intractable due to the number of computations. Therefore, a random subset of the samples generated by PPP was plotted. The size of this subset was chosen to be the closest integer of the larger set raised to the power of 0.90.90.9. Hence, the number of representations plotted monotonically increased with the number of PPP samples, yet remained computationally tractable for very large numbers of representations. This approach was used in all cases.

### v3

@@ B.3 Dependence on Latent Layer of Study
Figure 14: This network shows the PPP method applied over networks of $3$ hidden layer block, $18$ neurons and using input-output normalisation as described in Apps. A and D. PPP was applied to the $0^{\text{th}}$ hidden layer. The top five rows consist of the standard-tanh autoencoders, whereas the bottom five rows utilise isotropic-tanh. Everything else is equal between plots. Training proceeds rightwards over columns, in identical fashion to Fig. 1. One can observe that representations are absent in anisotropic rows 1, 2, and 4, likely a result of anti-alignment, whereas some axis-aligned structure is present in rows 3 and 5. Isotropic rows demonstrate larger, more continuous representational structures, but also demonstrate clustering such that they are not distributed smoothly over the representation space.
@@
Overall, in many networks tested, anisotropic structure appeared progressively clearer and denser in later latent layers, whilst isotropic networks tended to produce progressively more continuous distributions in later latent layers. This trend was observed across many of the networks studied, but sporadic differences between networks also occurred. This may be due to the increased number of an/isotropies in the forward pass or indirectly through the backwards pass, yet further study may elucidate the exact reason why progressively clearer structure may emerge in later layers.
@@ B.4 Results for Leaky-ReLU
Figure 18: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 33 hidden layer blocks and width of $18$ neurons per block. The top five rows consist of the autoencoder using the standard Leaky-ReLU, while the bottom five rows consist of the same autoencoders using the isotropic Leaky-ReLU. All five of the top five rows demonstrate clear anti-aligned quantisation in representations, predominantly in a single anti-aligned axis. These emerge early in training and become progressively more (approximately) discrete. This contrasts with the isotropic Leaky-ReLU examples, which approach a more continuous, evenly distributed embedding distribution, consistent with previous observations. Additionally, at relatively smaller magnitudes, the Leaky-ReLU results consistently show some deviation from the discrete-like rays; this remains unexplained.
Figure 19: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 22 hidden layer blocks and width of $18$ neurons per block. The top five rows consist of the autoencoder using the standard Leaky-ReLU, while the bottom five rows consist of the same autoencoders using the isotropic Leaky-ReLU. These results are highly consistent with Fig. 18, yet the 3rd row also shows the addition of axis-aligned structure.
@@
In particular, the $0^{\text{th}}$ layer standard Leaky-ReLU results always produced two narrow beams arranged antipodally, along a single diagonal. This anti-aligned structure occurs in most cases, and the standard Leaky-ReLU only rarely displays an axis-aligned structure. This single-diagonal may be indicative of the $S_{n}$ group, which does not include sign flips, unlike tanh’s $B_{n}$ group. Therefore, only a single diagonal may appear populated, as opposed to other diagonal directions, which could result from the various mirrors applicable in the $B_{n}$ group. This nuanced difference between demonstrations appears to closely follow from the differences between symmetry groups. These differing representations can be considered to support the influence of different symmetry groups on representations, with nuances following in a taxonomic fashion from the group structures. Particularly, the orthants containing $\pm\vec{1}$ are the only non-degenerate orthants produced by the $S_{n}$ symmetry, and the representations appear to be aligning with these, suggesting the remaining orthants appear not to produce computationally useful maps.
@@
B.4.1 Hyperoctahedral Leaky-ReLU
The prior results for Leaky-ReLU compared the contemporary standard permutation $S_{n}$ form of Leaky-ReLU against an isotropic form exhibiting $\mathrm{O}\left(n\right)$ equivariance. These results are now extended by considering hyperoctahedral, $B_{n}$, forms of Leaky-ReLU. The hyperoctahedral group is situated between these groups: $S_{n}\subset B_{n}\subset\mathrm{O}\left(n\right)$. Two approaches were considered to produce a $B_{n}$ form of Leaky ReLU. The first is an elementwise construction shown in Eqn. 8.
$\mathbf{f}\left(\vec{x};\left\{\hat{e}_{i}\right\}_{\forall\hat{e}_{i}}\right)=\sum_{i=1}^{n}f\left(\vec{x}\cdot\hat{e}_{i}\right)\hat{e}_{i}$ (8)
The univariate map, $f$, can be given piecewise by Eqn. 9.
[…]
As an alternative, a max-norm formulation was constructed as shown in Eqn. 10.
$\mathbf{f}\left(\vec{x};\left\{\hat{e}_{i}\right\}_{\forall\hat{e}_{i}}\right)=\left\{\begin{matrix}\alpha\vec{x}&:&\left\|\vec{x}\right\|_{\infty}<\epsilon\\ \vec{x}+\frac{\epsilon(\alpha-1)\vec{x}}{\left\|\vec{x}\right\|_{\infty}}&:&\left\|\vec{x}\right\|_{\infty}\geq\epsilon\end{matrix}\right$. (10)
Both formulations exhibit a hyperoctahedral symmetry about the standard basis. The results are shown below, including normalised and unnormalised cases, all performed on the CIFAR dataset.
Figure 22: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 11 hidden layer blocks and a width of $18$ neurons per block, in this example, normalisation was not used. The five rows consist of the autoencoder using the Hyperoctahedral Leaky-ReLU from Eqn. 8. All five rows show dense beams of discretised-like clusters of representations. Many of these are situated along two orthogonal standard basis directions. Additionally, beams are observed at angles of approximately 25 degrees away from the standard basis directions in several of the plots. The beams also become fainter through training, as they are likely distributed outside of the PPP threshold.
Figure 23: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 11 hidden layer blocks and a width of $18$ neurons per block, in this example, normalisation was not used. The five rows consist of the autoencoder using the Hyperoctahedral Leaky-ReLU from Eqn. 10. These are the same conditions as Fig. 22 but with the alternative implementation for hyperoctahedral Leaky-ReLU. In this plot, the top three rows show perpendicular beams aligning with the standard basis, whilst the bottom two rows show a single dense beam aligned with just one basis vector.
Both Figs. 22 and 23, demonstrate that the hyperoctahedral form of Leaky-ReLU also produces discrete-like distributions of activations. Additionally, these results are taken from the $0^{\text{th}}$ layer, indicating this is a non-trivial consequence. Normalisation was not utilised in these plots, and they are observationally similar to other permutation and hyperoctahedral plots without normalisation. This suggests that the algebraic symmetry of the functional form remains a useful predictor of biases that emerge in representations as a result of these functional form choices. The following plots of Figs. 24 and 25 demonstrate these hyperoctahedral variants for normalised cases.
Figure 24: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 22 hidden layer blocks and a width of $18$ neurons per block, in this example, normalisation was used. The five rows consist of the autoencoder using the Hyperoctahedral Leaky-ReLU from Eqn. 8. Both rows 11 and $5$ indicate a tendency for representation overdensities aligning with all standard basis vectors; in the remaining rows, a pattern is not so evident, but a faint alignment remains visible, particularly in rows 22 and $3$.
Figure 25: Demonstrates the same experiment as Fig. 24; however, with the hyperoctahedral formulation given in Eqn. 10. These results show a much stronger basis alignment in all cases, with a falloff around these more aligned beams. In all rows except row 11, there appear to be more dense representations in two antipodal directions, and then a slightly weaker density in the antipodal directions orthogonal to these.
Both Figs. 24 and 25 show that discretised representations continue to emerge in hyperoctahedral Leaky-ReLU networks with normalisation present too. Moreover, Fig. 25 shows qualitatively similar results to many hyperoctahedral Tanh networks of Figs. 9 and 10 as opposed to permutation Leaky-ReLU in Fig. 19. The latter is the same experiment as Fig. 25 but using permutation Leaky-ReLU. Overall, this strongly reinforces that maximal algebraic symmetries are a useful predictor, since the induced representational structure of Fig. 25, more closely aligns with other hyperoctahedral experiments as opposed to other Leaky-ReLU experiments, despite being a hyperoctahedral Leaky-ReLU analogue. This arrangement was observed in results for both Eqns. 8 and 10, but more often the latter.
A selection of further particularly indicative results is shown below, drawn from a variety of experiments.
Figure 26: The results of the PPP method applied to the $3^{\text{rd}}$ latent layer of an autoencoder with 33 hidden layer blocks and a width of $18$ neurons per block, in this example, normalisation was used. The five rows consist of the autoencoder using the Hyperoctahedral Leaky-ReLU from Eqn. 8. These results continue to show the pattern of representation overdensities situated over the standard basis directions.
Figure 27: The results of the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 33 hidden layer blocks and a width of $18$ neurons per block, in this example, normalisation was used. The five rows consist of the autoencoder using the Hyperoctahedral Leaky-ReLU from Eqn. 10. These results appear similar to those of Fig. 25, with dense beams of the standard basis vectors with some falloff.
Overall, these results continue to confirm that activation functions can induce unintended structure into distributions. They additionally show that the maximal algebraic symmetries continue to be a good predictor of the presence of this structure, with all results showing a clear trend. Moreover, the hyperoctahedral Leaky-ReLU ($B_{n}$-LeakyReLU) results are largely different from those of the standard permutation Leaky-ReLU ($S_{n}$-LeakyReLU) over the standard basis; yet, some similarities emerge in comparison to the hyperoctahedral Tanh. This indicates that these maximal symmetries may be a helpful way to categorise biases emerging from these primitives. Together, these are considered to strengthen the paper’s conclusions.
@@ B.5 MNIST Results
An initial example is displayed in Fig. 28 and 29 — the latter showing the same plot, with colour levels adjusted to make faint structure more observable for analysis. Both are included for comparability and transparency.
Figure 28: Displays the PPP method applied to the $0^{\text{th}}$ latent layer of an autoencoder with 33-hidden-layer blocks consisting of $18$ neurons per block. This architecture is comparable to those in App. D in every way except for the dimensions of CIFAR, which is exchanged for MNIST shape on input-output layers. Overall, one can observe that isotropic networks in the bottom five rows continue to produce smooth distributions that fill the representation space. In comparison, anisotropic networks on the top five rows tend to have many representations centred around the origin, and some produce axis-aligned features.
@@
Several of the $0^{\text{th}}$ layer plots show a closer grouping of representations about the origin. This is believed to be due to many of the MNIST examples’ components reaching the extreme values for the space, 0 and $1$, with fewer intermediate values. Often, this occurs repeatedly in specific patches due to the overlapping morphologies of MNIST numbers, which may produce clustered representations inherent to the dataset. This makes the MNIST dataset more anisotropic, featuring already dense, discrete clusters about the corners of an embedded hyper-cube — significantly different from the more intermediate RGB values arising from CIFAR. Slight band-groupings of representations are also sometimes observable, likely due to the same representation being projected in multiple planes, as well as the precision of the dataset, which produces slight clusterings. However, MNIST’s greater inherent anisotropy appears not to significantly alter the results from CIFAR, likely due to the layer bottleneck embedding these higher-dimensional representations more evenly in the latent layer during initialisation.
@@
Figure 30: Displays the PPP method applied to the $1^{\text{st}}$ latent layer of an autoencoder with 33-hidden-layer blocks consisting of $18$ neurons per block. Row-wise, these are the exact same networks as displayed in Fig. 28 and 29, but for the next latent layer. One can observe an axis-aligned structure in the top five rows, corresponding to anisotropic standard-tanh, and this is consistent with all previous experiments. In contrast, a smoother, approximately continuous distribution is observed for the bottom five rows, which are isotropic, and is consistent with this trend.
@@ Appendix C Analysis of Performance and Pathological Indications
Figure 32: These plots show the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green) reconstruction errors on test-set CIFAR against the number of training epochs. The networks have a width of $18$ neurons, and their depth increases to the right. In the leftmost plot, unnormalised isotropy has the lowest error by a very slight margin. The central two plots show that normalised isotropic networks have the lowest error by a large margin. The rightmost plot again shows unnormalised isotropy with the lowest error and a larger margin. Overall, a trend is not apparent, but isotropy tends to outperform by a greater margin.
@@
An instance of $32$ neurons was also computed, and this example did show unnormalised-anisotropic functions performing better as shown in Fig. 33. However, these were not the best-performing models overall.
Figure 33: This plot shows the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green), reconstruction error on test-set CIFAR against the number of training epochs. The networks have a width of $32$ neurons, and $2$ hidden layer blocks. In this counterexample, the unnormalised-anisotropic network had the lowest reconstruction error.
@@
Figure 34: This plot shows the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) reconstruction errors for test-set MNIST against the number of training epochs, network width increases rightward over columns, and depth increases downward over rows (flipped ordering from Fig. 31). It can be observed that reconstruction error is lower in wider and less-deep networks. Isotropic networks outperform anisotropic networks in every case, except for the 33-hidden-layer block and $32$ neurons, where anisotropic networks exhibit a slight marginal success. The best performing network is the isotropic network with 0 hidden layer blocks and $32$ neurons — this network is also thought to have the greatest representation capacity.
@@
Figure 35: These plots show the tanh-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green), reconstruction errors on test-set MNIST against the number of training epochs. The networks have a width of $18$ neurons, and a depth of 0 hidden layer blocks. It can be observed that all networks tend to a similar reconstruction error, except for the normalised-isotropy graph, which reaches a significantly lower reconstruction error.
@@
Figure 36: These plots show the Leaky-ReLU-like normalised anisotropic (orange) and normalised isotropic (blue) overlayed over the unnormalised anisotropic (red) and unnormalised isotropic (green) reconstruction errors on test-set CIFAR against the number of training epochs. The networks have a width of $18$ neurons, and their depth increases to the right. In all instances, normalised or unnormalised isotropic-Leaky-ReLU reaches the lowest reconstruction error by a large margin. Anisotropic networks also appear to exhibit significantly greater noise and sample-to-sample variation in test-set accuracies across the differing training epochs.
@@ Appendix D Further Details: Network Architectures and Training
This depicts the architectures used across all results, where $n$ varies in width and $L$ varies in depth of the autoencoder model. The latent layer of study is indexed by integers $k\in\left[0,L\right]$. For MNIST LeCun et al. [2010] models, the tensor representations are changed from $\left[32,32,1\right]\rightarrow\left[28,28,1\right]$ and $\left[3072\right]\rightarrow\left[784\right]$ in all instances, with corresponding changes in affine maps. This ensures these results follow from a similar and comparable model.
Finally, following from App. A, the MNIST and CIFAR datasets are sometimes normalised over their training sets to reduce statistical geometry artefacts. This involves an element-wise calculation of the mean tensor over all training samples, as well as for the standard deviation statistic. Where the standard deviation is found to be 0, those elements are replaced with the value $1$. When normalisation is used, every ingoing sample per batch is standardised by subtracting the mean and dividing by the standard deviation, where the non-zero standard deviation is made equal to one.
Throughout all experiments, a batch size of $24$ was maintained, alongside a learning rate for PyTorch’s Paszke et al. [2019] momentum gradient descent, with momentum $0.9$, a learning rate $0.08$, and using a mean-squared error loss for reconstruction. Parameter initialisation was the standard Xavier-Normal Glorot and Bengio [2010] for weights and zero-initialised for biases, this is due to their left-, right- probabilistic invariances Bird [2025a]. However, standard multivariate normal or orthogonal initialisation would have been equally valid choices. Still, Xavier-Normal was deemed more conventional, though its gain is only optimised for anisotropic activation functions Bird [2025a]. All values were kept in float64.
Networks were trained for a total of $125$ epochs; however, intermediate measurements were gathered after training for each of the incremental following amount of epochs: $\left[1,1,2,2,2,5,5,6,6,10,10,10,15,15,15,20\right]$ corresponding to a cumulative epoch total of $\left[1,2,4,6,8,13,18,24,30,40,50,60,75,90,105,125\right]$. This was chosen arbitrarily, but was found to give meaningful insight into the formation of the representation structures.
Data was gathered using combination-PPP, with $\epsilon=0.75$, which was found to be a good overall value. Plots were compiled using Gaussians centred at each representation’s projected coordinates, and the colour map indicates representation density per region, with a maximum value of the mean plus five standard deviations and a minimum of zero. Additionally, in the anisotropic cases, so many representations became aligned with the axes that computing the Gaussian maps became intractable due to the number of computations. Therefore, a random subset of the samples generated by PPP was plotted. The size of this subset was chosen to be the closest integer of the larger set raised to the power of $0.9$. Hence, the number of representations plotted monotonically increased with the number of PPP samples, yet remained computationally tractable for very large numbers of representations. This approach was used in all cases.